# Rewarding the Journey, Not Just the Destination: A Composite Path and Answer Self-Scoring Reward Mechanism for Test-Time Reinforcement Learning

### v3

@@ 1 Introduction
Reinforcement Learning (RL) (Kaelbling et al., 1996; Dong et al., 2024) has emerged as a powerful paradigm for advancing the capabilities of pre-trained Large Language Models (LLMs), driving significant progress in complex reasoning domains, e.g., mathematics Setlur et al. (2024); Gao et al. (2024); Albright & Andemicael (2025) and code generation (Guo et al., 2025; Achiam et al., 2023; Team, 2025; Wang et al., 2024; Islam et al., 2024). However, current RL approaches predominantly rely on explicit external supervision through ground-truth labels (Luong et al., 2024; Ouyang et al., 2022; Shao et al., 2024) or human preference data to construct reward functions. This dependency creates a fundamental scalability bottleneck: as tasks grow in complexity and volume, large-scale annotation becomes increasingly impractical, hindering the continuous evolution of state-of-the-art models.
@@
This limitation naturally motivates an alternative paradigm where LLMs autonomously improve through RL on unlabeled data, learning directly from continuous experience streams (Zhang et al., 2025c; Xiong et al., 2025; Huang et al., 2024). However, the core challenge in this paradigm lies in reward estimation during inference without access to ground-truth information. Test-Time Reinforcement Learning (TTRL) (Zuo et al., 2025) formalizes this setting by enabling parameter updates using unlabeled test data—a promising direction that has recently gained significant attraction. TTRL addresses the reward challenge by sampling multiple responses per problem and constructing pseudo-labels through self-consistency consensus via majority voting. The approach relies on the model’s intrinsic confidence as a proxy metric when external rewards are unavailable: answers consistently reproduced across multiple trials are considered higher-confidence and thus more likely to be correct. This mechanism mirrors human problem-solving, where conclusions verified through diverse methods strengthen our confidence in their validity.
@@
While using confidence as a correctness proxy aligns with cognitive principles, a critical question remains: how is confidence actually manifested in the reasoning process of LLMs? TTRL(Zuo et al., 2025) adopts self-consistency via majority voting as its confidence proxy. However, this choice has clear limitations. When initial pseudo-labels derived from voting are biased or incorrect, reflecting flaws in the pre-trained model’s prior knowledge, the model risks reinforcing erroneous consensus. This raises an important motivation for our work: can we exploit other forms of prior knowledge within LLMs to mitigate these limitations and obtain more reliable self-reward signals? Since LLMs fundamentally operate as next-token predictors, their most direct internal state evidence resides in the probability distribution over candidate tokens. Different measures of this distribution naturally yield alternative notions of model confidence. For instance, the entropy of the distribution reflects uncertainty—lower entropy indicates higher confidence. Similarly, the probability assigned to the top-1 token represents the model’s certainty in its preferred choice. Moreover, the margin between the top-1 and top-2 token probabilities captures decisiveness; a larger margin implies the model is less hesitant, and thus more confident.
@@ 2.1 RL for Reasoning
RL plays a critical role in enhancing the instruction-following and reasoning capabilities of LLMs (Guo et al., 2025; Ouyang et al., 2022). Over time, research in this direction has evolved through three main paradigms, each progressively reducing reliance on external supervision. Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Zheng et al., 2023; Dai et al., 2023), aligns base models with human preferences using learnable reward models trained on annotated preference data. Methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) are widely employed, but the dependence on large-scale human feedback limits scalability. Reinforcement Learning with Verifiable Rewards (RLVR) (Luong et al., 2024; Shao et al., 2024), replaces preference-based supervision with rule-based reward functions grounded in gold-standard answers. This class of methods, which has been particularly effective for math and code generation tasks, reduces the need for nuanced human feedback but still requires labeled datasets to provide ground-truth answers. Reinforcement Learning from Internal Feedback (RLIF) (Zhang et al., 2025a; Zuo et al., 2025; Zhao et al., 2025), eliminates explicit supervision altogether by leveraging intrinsic signals from LLMs themselves. The underlying idea is that LLMs already possess a substantial amount of knowledge, which—if effectively organized—can support correct reasoning. However, direct inference after supervised training often fails to fully exploit this potential, requiring reinforcement learning to discover strategies that better coordinate and utilize existing knowledge. Conceptually, RLIF aims to formalize this intuition by framing the organization and utilization of pre-existing knowledge within LLMs as a policy optimization problem, wherein intrinsic feedback signals serve as implicit rewards guiding the emergence of more effective reasoning strategies. For example, EMPO (Zhang et al., 2025a) incentivizes reasoning by minimizing entropy in a latent semantic space, while TTRL (Zuo et al., 2025) generates pseudo-labels via self-consistency consensus among multiple sampled responses and uses them to compute rewards. These approaches mark a shift toward self-rewarding reinforcement learning, aiming to enable LLMs to improve directly from unlabeled data. Our work also falls within the scope of RLIF.
@@ 2.2 Confidence-Based Reward
In fully unsupervised settings (Zhang et al., 2025b; Wei et al., 2025), where ground-truth labels are unavailable, directly optimizing for the correctness of Large Language Model (LLM) outputs is infeasible. A natural alternative is to exploit intrinsic proxy metrics that correlate strongly with correctness. Among these, model confidence (Xiong et al., 2023; Tripathi et al., 2025; Tian et al., 2025) has emerged as a particularly informative indicator, mirroring human cognition, in which confidence serves as an internal estimate of reliability in the absence of external supervision. Current confidence-based reward methods have evolved along three complementary trajectories. One paradigm quantifies confidence through the log-likelihood of generated sequences, rewarding responses that the model assigns higher probability; for instance, RLSC (Li et al., 2025) reinforces generation by maximizing sequence log-likelihood. Another paradigm centers on entropy minimization, where lower entropy implies greater certainty—an idea instantiated by RENT (Prabhudesai et al., ), which encourages more decisive and coherent outputs. A third direction evaluates self-consistency across multiple samples, as in TTRL (Zuo et al., 2025), which derives pseudo-labels via consensus among sampled responses and reinforces outputs consistent with the emergent majority. While these confidence-based methods have yielded notable gains in reasoning and stability, they predominantly rely on a single dimension of confidence—whether likelihood, entropy, or consensus. This leaves open an important question: how can multiple complementary confidence signals be integrated to construct more reliable and fine-grained self-reward mechanisms? Our work addresses this gap by proposing a unified reward framework that leverages diverse intrinsic feedback sources to guide more robust self-improvement in LLMs.
@@ 2.3 Test-Time Adaptation
Test-Time Adaptation (TTA) (Sun et al., 2017; Maria Carlucci et al., 2017; Schneider et al., 2020) refers to updating a model on unlabeled test data during inference, aiming to mitigate performance degradation caused by distribution shifts between training and testing environments. Early studies in computer vision primarily explored unsupervised objectives that promote robustness to domain shifts. More recently, the TTA paradigm has been extended to LLMs and reasoning tasks. For instance, Tent (Wang et al., 2020) adapts models by minimizing the entropy of predictions, under the assumption that well-adapted models should produce certain (i.e., low-entropy) outputs on test samples. Similarly, TTRL (Zuo et al., 2025) introduces test-time reinforcement learning, where rewards are derived from self-consistency via majority voting among sampled responses. Despite these advances, most existing TTA approaches remain limited in scope: they either focus on generic uncertainty reduction or rely on outcome-level consensus, offering only indirect supervision over the underlying reasoning process. This limitation highlights the need for more expressive test-time reward mechanisms that can jointly assess both answer reliability and reasoning quality, thereby enabling stronger and more principled adaptation in LLMs.
@@ 3.1 Dual-Calibration Answer Reward (DCAR)
To further enhance the reliability of our confidence-calibrated pseudo-label $y^{*}$, we introduce a credibility score $\mathcal{S}_{cred}$ that assesses consensus quality relative to the most confident response, effectively implementing soft curriculum learning. This approach is grounded in a key hypothesis: a consensus derived from high-confidence responses is more reliable than one based on diverse low-confidence outputs. Building on this principle, we define the credibility metric using two fundamental concepts: General Group $\mathcal{C}_{\text{General}}$ and Elite Response $\mathcal{C}_{\text{Elite}}$. Specifically, the General Group contains all responses that agree with the pseudo-label $y^{*}$. The group’s confidence $\mathcal{C}_{\text{General}}$ is defined as the maximum confidence within this subset, representing the strongest supporting evidence for the consensus, and the $\mathcal{C}_{\text{Elite}}$ is the response among all NN candidates with the highest confidence:
@@ 4.1 Experimental Setup
Benchmarks. To comprehensively evaluate COMPASS’s generalizability, we employ diverse backbone models spanning different scales and specializations: LLaMA-3.2-1B-Instruct (Meta, 2024) as an instruct-tuned model, Qwen2.5-Math-1.5B (Yang et al., 2025) as a mathematically specialized base model, and Qwen2.5-7B (Yang et al., 2025) as a general-purpose base model. We conduct evaluation on GPQA-Diamond (Rein et al., 2024), a challenging subset of the Graduate-Level Google-Proof Question Answering benchmark, alongside three mathematical reasoning benchmarks: AIME 2024 (Yuan et al., 2024), AMC (Yuan et al., 2024), and MATH-500 (Hendrycks et al., 2021). As TTRL (Zuo et al., 2025) pioneered the test-time reinforcement learning paradigm, we primarily compare COMPASS against both backbone models and TTRL baselines to validate its effectiveness in enabling autonomous self-improvement.
Metrics. We evaluate COMPASS independently on each benchmark using the pass@k protocol following DeepSeek-R1 (Guo et al., 2025), reporting pass@1 scores with non-zero temperature sampling. Specifically, we generate 16 responses per question (4 for 32k context length) using temperature 0.6 and top-pp 0.95. The pass@1 score is computed as: $\text{pass@1}=\frac{1}{k}\sum_{i=1}^{k}p_{i},$, where $p_{i}\in{0,1}$ indicates the correctness of the $i$-th response.
Implementation Details. We implement COMPASS using GRPO (Shao et al., 2024) independently on each benchmark. For optimization, we employ AdamW with a cosine learning rate schedule (peak $5\times 10^{-7}$). During rollout, we sample 64 responses (temperature 0.6, except 1.0 for Qwen2.5-Math) for pseudo-label estimation via voting, then downsample to 32 responses per prompt for training. This design maintains computational efficiency while achieving strong performance. For models with fewer than 7B parameters, we follow TTRL settings with 10, 30, and 80 episodes for MATH-500, AMC, and AIME 2024 respectively, scaled by dataset size. For Qwen2.5-7B, we reduce epochs to 2, 8, and 20 (approximately 20% of TTRL’s budget) to evaluate efficiency in computationally constrained regimes.

### v4

@@ 1 Introduction
Reinforcement Learning (RL) Kaelbling et al. (1996); Dong et al. (2024) has emerged as a powerful paradigm for advancing the capabilities of pre-trained Large Language Models (LLMs), driving significant progress in complex reasoning domains, e.g., mathematics Setlur et al. (2024); Gao et al. (2024); Albright & Andemicael (2025) and code generation Guo et al. (2025); Achiam et al. (2023); Team (2025); Wang et al. (2024); Islam et al. (2024). However, current RL approaches predominantly rely on explicit external supervision through ground-truth labels Luong et al. (2024); Ouyang et al. (2022); Shao et al. (2024) or human preference data to construct reward functions. This dependency creates a fundamental scalability bottleneck: as tasks grow in complexity and volume, large-scale annotation becomes increasingly impractical, hindering the continuous evolution of state-of-the-art models.
@@
This limitation naturally motivates an alternative paradigm where LLMs autonomously improve through RL on unlabeled data, learning directly from continuous experience streams Zhang et al. (2025c); Xiong et al. (2025); Huang et al. (2024). However, the core challenge in this paradigm lies in reward estimation during inference without access to ground-truth information. Test-Time Reinforcement Learning (TTRL) Zuo et al. (2025) formalizes this setting by enabling parameter updates using unlabeled test data—a promising direction that has recently gained significant attraction. TTRL addresses the reward challenge by sampling multiple responses per problem and constructing pseudo-labels through self-consistency consensus via majority voting. The approach relies on the model’s intrinsic confidence as a proxy metric when external rewards are unavailable: answers consistently reproduced across multiple trials are considered higher-confidence and thus more likely to be correct. This mechanism mirrors human problem-solving, where conclusions verified through diverse methods strengthen our confidence in their validity.
@@
While using confidence as a correctness proxy aligns with cognitive principles, a critical question remains: how is confidence actually manifested in the reasoning process of LLMs? TTRLZuo et al. (2025) adopts self-consistency via majority voting as its confidence proxy. However, this choice has clear limitations. When initial pseudo-labels derived from voting are biased or incorrect, reflecting flaws in the pre-trained model’s prior knowledge, the model risks reinforcing erroneous consensus. This raises an important motivation for our work: can we exploit other forms of prior knowledge within LLMs to mitigate these limitations and obtain more reliable self-reward signals? Since LLMs fundamentally operate as next-token predictors, their most direct internal state evidence resides in the probability distribution over candidate tokens. Different measures of this distribution naturally yield alternative notions of model confidence. For instance, the entropy of the distribution reflects uncertainty—lower entropy indicates higher confidence. Similarly, the probability assigned to the top-1 token represents the model’s certainty in its preferred choice. Moreover, the margin between the top-1 and top-2 token probabilities captures decisiveness; a larger margin implies the model is less hesitant, and thus more confident.
@@ 2.1 RL for Reasoning
RL plays a critical role in enhancing the instruction-following and reasoning capabilities of LLMs Guo et al. (2025); Ouyang et al. (2022). Over time, research in this direction has evolved through three main paradigms, each progressively reducing reliance on external supervision. Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Zheng et al. (2023); Dai et al. (2023), aligns base models with human preferences using learnable reward models trained on annotated preference data. Methods such as Proximal Policy Optimization (PPO) Schulman et al. (2017) are widely employed, but the dependence on large-scale human feedback limits scalability. Reinforcement Learning with Verifiable Rewards (RLVR) Luong et al. (2024); Shao et al. (2024), replaces preference-based supervision with rule-based reward functions grounded in gold-standard answers. This class of methods, which has been particularly effective for math and code generation tasks, reduces the need for nuanced human feedback but still requires labeled datasets to provide ground-truth answers. Reinforcement Learning from Internal Feedback (RLIF) Zhang et al. (2025a); Zuo et al. (2025); Zhao et al. (2025), eliminates explicit supervision altogether by leveraging intrinsic signals from LLMs themselves. The underlying idea is that LLMs already possess a substantial amount of knowledge, which—if effectively organized—can support correct reasoning. However, direct inference after supervised training often fails to fully exploit this potential, requiring reinforcement learning to discover strategies that better coordinate and utilize existing knowledge. Conceptually, RLIF aims to formalize this intuition by framing the organization and utilization of pre-existing knowledge within LLMs as a policy optimization problem, wherein intrinsic feedback signals serve as implicit rewards guiding the emergence of more effective reasoning strategies. For example, EMPO Zhang et al. (2025a) incentivizes reasoning by minimizing entropy in a latent semantic space, while TTRL Zuo et al. (2025) generates pseudo-labels via self-consistency consensus among multiple sampled responses and uses them to compute rewards. These approaches mark a shift toward self-rewarding reinforcement learning, aiming to enable LLMs to improve directly from unlabeled data. Our work also falls within the scope of RLIF.
@@ 2.2 Confidence-Based Reward
In fully unsupervised settings Zhang et al. (2025b); Wei et al. (2025), where ground-truth labels are unavailable, directly optimizing for the correctness of Large Language Model (LLM) outputs is infeasible. A natural alternative is to exploit intrinsic proxy metrics that correlate strongly with correctness. Among these, model confidence Xiong et al. (2023); Tripathi et al. (2025); Tian et al. (2025) has emerged as a particularly informative indicator, mirroring human cognition, in which confidence serves as an internal estimate of reliability in the absence of external supervision. Current confidence-based reward methods have evolved along three complementary trajectories. One paradigm quantifies confidence through the log-likelihood of generated sequences, rewarding responses that the model assigns higher probability; for instance, RLSC Li et al. (2025) reinforces generation by maximizing sequence log-likelihood. Another paradigm centers on entropy minimization, where lower entropy implies greater certainty—an idea instantiated by RENT Prabhudesai et al. , which encourages more decisive and coherent outputs. A third direction evaluates self-consistency across multiple samples, as in TTRL Zuo et al. (2025), which derives pseudo-labels via consensus among sampled responses and reinforces outputs consistent with the emergent majority. While these confidence-based methods have yielded notable gains in reasoning and stability, they predominantly rely on a single dimension of confidence—whether likelihood, entropy, or consensus. This leaves open an important question: how can multiple complementary confidence signals be integrated to construct more reliable and fine-grained self-reward mechanisms? Our work addresses this gap by proposing a unified reward framework that leverages diverse intrinsic feedback sources to guide more robust self-improvement in LLMs.
@@ 2.3 Test-Time Adaptation
Test-Time Adaptation (TTA) Sun et al. (2017); Maria Carlucci et al. (2017); Schneider et al. (2020) refers to updating a model on unlabeled test data during inference, aiming to mitigate performance degradation caused by distribution shifts between training and testing environments. Early studies in computer vision primarily explored unsupervised objectives that promote robustness to domain shifts. More recently, the TTA paradigm has been extended to LLMs and reasoning tasks. For instance, Tent Wang et al. (2020) adapts models by minimizing the entropy of predictions, under the assumption that well-adapted models should produce certain (i.e., low-entropy) outputs on test samples. Similarly, TTRL Zuo et al. (2025) introduces test-time reinforcement learning, where rewards are derived from self-consistency via majority voting among sampled responses. Despite these advances, most existing TTA approaches remain limited in scope: they either focus on generic uncertainty reduction or rely on outcome-level consensus, offering only indirect supervision over the underlying reasoning process. This limitation highlights the need for more expressive test-time reward mechanisms that can jointly assess both answer reliability and reasoning quality, thereby enabling stronger and more principled adaptation in LLMs.
@@ 3.1 Dual-Calibration Answer Reward (DCAR)
To further enhance the reliability of our confidence-calibrated pseudo-label $y^{*}$, we introduce a credibility score $\mathcal{S}_{cred}$ that assesses consensus quality relative to the most confident response, effectively implementing soft curriculum learning. This approach is grounded in a key hypothesis: a consensus derived from high-confidence responses is more reliable than one based on diverse low-confidence outputs. Building on this principle, we define the credibility metric using two fundamental concepts: General Group $\mathcal{C}_{\text{General}}$ and Elite Response $\mathcal{C}_{\text{Elite}}$. Specifically, the General Group contains all responses that agree with the pseudo-label $y^{*}$. The group’s confidence $\mathcal{C}_{\text{General}}$ is defined as the maximum confidence within this subset, representing the strongest supporting evidence for the consensus, and the $\mathcal{C}_{\text{Elite}}$ is the response among all $N$ candidates with the highest confidence:
@@ 4.1 Experimental Setup
Benchmarks. To comprehensively evaluate COMPASS’s generalizability, we employ diverse backbone models spanning different scales and specializations: LLaMA-3.2-1B-Instruct Meta (2024) as an instruct-tuned model, Qwen2.5-Math-1.5B Yang et al. (2025) as a mathematically specialized base model, and Qwen2.5-7B Yang et al. (2025) as a general-purpose base model. We conduct evaluation on GPQA-Diamond Rein et al. (2024), a challenging subset of the Graduate-Level Google-Proof Question Answering benchmark, alongside three mathematical reasoning benchmarks: AIME 2024 Yuan et al. (2024), AMC Yuan et al. (2024), and MATH-500 Hendrycks et al. (2021). As TTRL Zuo et al. (2025) pioneered the test-time reinforcement learning paradigm, we primarily compare COMPASS against both backbone models and TTRL baselines to validate its effectiveness in enabling autonomous self-improvement.
Metrics. We evaluate COMPASS independently on each benchmark using the pass@k protocol following DeepSeek-R1 Guo et al. (2025), reporting pass@1 scores with non-zero temperature sampling. Specifically, we generate 16 responses per question (4 for 32k context length) using temperature 0.6 and top-$p$ 0.95. The pass@1 score is computed as: $\text{pass@1}=\frac{1}{k}\sum_{i=1}^{k}p_{i},$, where $p_{i}\in{0,1}$ indicates the correctness of the $i$-th response.
Implementation Details. We implement COMPASS using GRPO Shao et al. (2024) independently on each benchmark. For optimization, we employ AdamW with a cosine learning rate schedule (peak $5\times 10^{-7}$). During rollout, we sample 64 responses (temperature 0.6, except 1.0 for Qwen2.5-Math) for pseudo-label estimation via voting, then downsample to 32 responses per prompt for training. This design maintains computational efficiency while achieving strong performance. For models with fewer than 7B parameters, we follow TTRL settings with 10, 30, and 80 episodes for MATH-500, AMC, and AIME 2024 respectively, scaled by dataset size. For Qwen2.5-7B, we reduce epochs to 2, 8, and 20 (approximately 20% of TTRL’s budget) to evaluate efficiency in computationally constrained regimes.